### v1

@@ 3.1. Design Goals
The central purpose of Bonsai is to realize a feed that is both explicit and personalized. We operationalize this purpose by pursuing three design goals detailed below: intent fidelity, transparency, and controllability.
@@
Explicitly stated intentions govern the sourcing, filtering, and ranking of content in the feed. Users should have an interface that allows them to specify directly what they want—and do not want—to see. This goal responds to prior work (§2.1) showing that optimizing for revealed preferences (e.g., clicks, dwell time) often drives regrettable or addictive use  (Tran et al., 2019a; Cho et al., 2021; Milli et al., 2025)). This allows users to isolate the personalization signal to what they explicitly instruct the system to do, preventing hidden optimization toward engagement and ensuring that their social media use aligns with their intended goals.
3.1.2. Controllability.
Users set and adjust feed goals without specialized vocabulary or interfaces. The interface between users and the feed algorithm should be simple, flexible, and expressive. Lowering the cost of articulation broadens who can personalize feeds and reduces drift toward opaque defaults. This design goal draws on HCI systems (§2.2) that show lightweight, accessible interventions are effective in helping people regulate their technology use (Hiniker et al., 2016; Grüning et al., 2023; Ko et al., 2015). Controllability is also important to ensure intent fidelity, as increasing the flexibility and ease with which users specify their intentions ensures alignment between intention and feed.
@@
The feed creation process is procedurally transparent. Users should clearly understand the process by which their feed is generated, and how different aspects of the configuration impact the resulting feed. This design goal builds on research (§2.3) showing that people frequently misunderstand how feeds operate and consequently miscalibrate their trust or behavior (Eslami et al., 2016; Rader and Gray, 2015), as well as on scholarship critiquing “black-box” notions of transparency and urging attention to the processes through which systems operate (Ananny and Crawford, 2018). Transparency is essential to make controllability meaningful: only when users understand how their input shapes outputs can they refine and trust the system.
@@ 3.2. Interface Design and System Walkthrough
We designed Bonsai based on the aforementioned landscape analysis and design goals. During prototyping, we iteratively refined defaults and interaction flows in response to formative feedback from researchers involved in the project and informal demos to colleagues. Two key resulting changes from this iterative process were: 1) the creation of an initial prompt interface (see §3.2.1) to bootstrap the feed creation process; and 2) the addition of different ranking algorithms that prioritize recency, popularity, or topical fit (see §3.2.4).
We implemented Bonsai for Bluesky for normative and pragmatic reasons. Bluesky is a microblogging platform that is open, decentralized, built on the AT protocol, and designed to give users more control over their experience, content, and data. The AT protocol, therefore, facilitates the easy implementation of a feed curation system, and Bluesky users are also more likely to be interested in such an application.
Figure 3 shows the resulting Bonsai interface. The user experience unfolds through four primary steps, which we describe below.
@@ 3.2.2. Selecting sources
Next, users specify the sources that will populate their feed (Fig. 3B). Based on the initial description, Bonsai suggests feeds, starter packs, lists, search queries, accounts, and hashtags. Users can accept, edit, or remove these suggestions, as well as manually add new sources. Sources can be accounts, other feeds, search terms, and ‘starter packs’ (curated sets of accounts associated with a specific topic).
@@ 3.2.5. Generating and using the feed
After completing these steps, the user can click “Generate Now” to initiate the feed generation process. A few seconds after initiating generation, Bonsai may suggest additional sources to the user, based on the edits made to the content preferences. After generation, the feed appears as a custom feed option in the user’s Bluesky app. Bonsai continues to update the feed in the background. If they wish to edit their feed, users can always make changes on the feed customization page. They can modify preferences, add new exclusions, adjust source selections, and even create new feeds.
@@ 3.3. System Architecture
We propose a generalizable system architecture for generating intentional and personalized social media feeds. This architecture comprises four components (Planning, Sourcing, Curation, and Ranking), depicted in Fig. 7. While our implementation targets Bluesky, the modular design is platform-agnostic: each component can be adapted to different ecosystems by plugging in platform-specific APIs, content databases, or models. We chose Bluesky because its open, decentralized AT Protocol makes external feed customization feasible, and its user community is predisposed to experimenting with novel personalization tools.
@@ 3.3.1. Planning
The Planning component translates a user’s natural language prompt into an initial feed configuration. This includes suggested sources (e.g., feeds, lists, hashtags) and content preferences (e.g., “include posts about” and “limit posts about”). Because this step involves semi-autonomous interpretation of natural language, we employ a language model (LLM) agent. In our Bluesky prototype, the agent queries an internal SQLite database of candidate sources (with previously scraped feeds, lists, and ‘starter packs’) and generates search queries and filter suggestions from the prompt. The resulting configuration is surfaced to the user for review and refinement. Our implementation used GPT-4o via Azure; the Planning stage can be instantiated with other models or retrieval systems, depending on the target platform. More details about the database construction are available in the Appendix B.
@@ 3.3.3. Curating
The Curating component applies the user’s stated preferences to the sourced posts and assigns each item a relevance score. In principle, this stage can be implemented with classifiers or rules tailored to the platform; in our case, we leveraged a multimodal LM (GPT-4o) to handle Bluesky’s mix of text and images and to align directly with users’ natural language prompts. Each post received a score on a 0–10 scale: “strongly prefer” (8–10), “prefer” (5–7), “unspecified” (3–4). “show less often” (1–2), ‘ ‘never show” (0). Posts scored as zero were filtered out. Although our prototype relied on a relatively heavyweight model, the curation logic could be reimplemented with lighter models, heuristics, or hybrid pipelines, depending on deployment constraints.
@@ 3.3.4. Ranking
The Ranking component orders the curated posts into a final feed according to user-specified priorities. We adopted three widely used dimensions from existing feed algorithms: relevance, recency, and popularity (Covington et al., 2016; Guess et al., 2023). Relevance scores came from the Curating stage; popularity was computed as a weighted engagement score (likes + 3 $\times$ reposts + 2 $\times$ replies); and recency was calculated via reverse-chronological order. To combine these, we applied a Weighted Borda Count, with weights set by the user’s chosen ranking style (e.g., “Focused,” “Fresh,” “Balanced,” “Trending”). This structure generalizes readily: other platforms may substitute alternative engagement metrics, quality signals, or normative objectives. We provide details about the Weighted Borda Count algorithm in the Appendix Alg. 1.
@@ 4. Methods
We conducted a field study with 15 participants who used Bonsai for a median of 11 days (min: 6 days; max: 25 days; avg: 12.27 days). We aimed for approximately one week of use for each participant, though the exact number of days varied based on availability. We interviewed participants at the beginning and end of the study in order to understand the details of their experience. The study lasted one month between July 29th and August 29th, 2025.
@@ 4.1. Participants
We recruited participants primarily through social media, creating a visual flyer that we shared on Bluesky and in relevant Slack groups. We also conducted targeted outreach on Bluesky to engage frequent users of the platform and to enlist non-academics in the study. From those who filled out the interest survey, we contacted participants we could validate as real people and as regular Bluesky users. In total, 15 participants enrolled in the study. Participants were paid \$50 for completing both interviews, and all procedures were approved by [Anonymous institution’s] IRB.
@@ 4.2. Procedure
The study comprised an onboarding interview, at least a week of using Bonsai feeds, and a follow-up interview. Both interviews were semi-structured and lasted 45 minutes. In the first interview, we explored participants’ general social media habits, asking what value they currently derive from feeds, what they feel is missing, and what they wish feeds would provide. We then walked them through the feed generation process in Bonsai, asking them to think aloud as they selected sources, wrote prompts, and configured ranking options. During the trial week, participants were encouraged to use Bonsai in their everyday browsing, creating and refining feeds as they wished. The second interview took a more reflective and evaluative form. We asked participants to describe how they attempted to convey their intent in the feed, whether they felt successful, and what challenges they encountered. We also probed the broader pros and cons of intentionality, natural language input, and feed-level control, as well as participants’ overall impressions of the system. In several cases, we revisited the feed creation process to observe how their strategies evolved. In both interviews, we examined the feeds participants had generated, discussing whether the content matched their expectations and probing moments when posts seemed out of place. We include the full interview protocols in the Appendix sections E and F.
@@ 4.3. Data Analysis
To analyze our qualitative interview data and identify main findings, we performed thematic analysis inspired by grounded theory (Braun and Clarke, 2006). Participant interviews were recorded and automatically transcribed on Zoom. We uploaded the anonymized transcripts to Atlas.ti, where they were further refined for accuracy. The first and second authors then engaged in open coding for two full interviews each, discussing and honing the initial codes with the broader research team. We iteratively expanded the codebook and continued to code interviews as they were being conducted. Interviews continued until thematic saturation was reached, except for one additional interview, which was included following saturation to capture more non-academic perspectives. To add richness to our findings, we also captured the Bonsai interaction logs and the natural language prompts from participants. The natural language prompts were tagged with specific user requests, such as a topical focus or a request to filter specific content. Additionally, we manually analyzed the logs for each participant to identify the actions they took on the platform between interviews, such as editing a feed or creating a new feed. These results are incorporated throughout the findings section.
@@ 5. Results
We discuss how participants used Bonsai during the study period, focusing on the prompts they authored, the actions they performed, and the ways they evaluated their feeds. We structure our findings around our three design goals: (1) how participants imprinted their intentions on social media feeds, (2) how they exercised control over explicit and personalized feeds, and (3) how they made sense of the procedural transparency of feed generation. Our 15 participants built between 1 and 4 feeds each for a total of 36 feeds. They manually re-generated their feeds between 2 and 16 times for a total of 103 times. Per week, they consumed Bonsai feeds between 8 and 264 times each for a total of 1413 requests.
@@ 5.1. Imprinting intentions on social media feeds
5.1.1. Participants specified topical interests using Bonsai, to create both general and specialized feeds.
@@
In the remaining 16 feeds, participants used Bonsai to build feeds for specific topics or hobbies they were interested in. P7 created a feed focused on cinematography (“I want to see film festival news, film critics, artists’ sharing their experiences or new releases, experimental / AI films”), P11 on college sports (“I want to see all posts related to US college basketball and college football”), and P15 “a videogame development feed.” Before using Bonsai, some participants had already used complementary topical feeds, such as P6, who followed a feed of political science journals, while P9 would repeatedly search for common keywords by clicking on the same keywords from their saved search history – effectively creating a pseudo-feed. For this participant, Bonsai offered a much easier way to create a topical feed suited to their interests.
@@
These desires for better filters stemmed from participants’ negative social media experience, particularly on Twitter. Some users sought to improve their mental health on platforms like X, which P8 called “toxic” and “unbearable,” P10 reported feeling “anxious when using it,” and P6 felt made them “deeply unhappy.” Others felt that engagement-based feeds simply were not effective at filtering out content, even when given explicit instructions. P8 said that on Twitter they “would always say, okay, I want to see less of that, and I wouldn’t see less of it,” while P15 held the same opinion of Bluesky, where “even though you tell it ‘show less of this’, it keeps displaying it.” P4 told the story of a user whose academic content they enjoyed, but who also posted frequently about football—a topic they had no interest in. In response, P4 tried to block football-relevant keywords, but found “with the existing methods, you can achieve certain results, they are certainly useful in some contexts, but they’re not very specific, they’re not very, sort of fine-grained.” P13 shares a similar experience, saying ‘on Bluesky, I would say [filtering] is terrible. I have set up 30 keywords to mute things, so that I don’t see political content on Bluesky, but it still manages to creep onto my feed.”
@@
Although participants felt intentional feeds could be liberating, some participants felt perfectly in control of their existing engagement-based feeds. P13 for example explained “on Twitter, […] I could just hit ’not interested’ and then I won’t see stuff,” and that they feel “I have a lot of self-control on what I click on to shape my algorithmic feed, and so because of that, I can get a really good feed on Twitter.” P9 similarly felt confident that platforms could infer their intent correctly: “once I give somebody a thumb [or a] heart, [the platform] can more quickly capture my interests and then start recommending more things similar to that”. Even those who felt more lukewarm about their ability to manage engagement-based feeds still exerted additional controls over their feeds by blocking and muting accounts and keywords to filter out content they did not wish to see.
@@
Through our design, participants cleverly used natural language to convey specific and nuanced desires. For example, participants used humor and specific cultural references to help the AI “situate” itself within the type of desired content. Nuanced references could allow participants to convey complex semantic ideas with just a few words. These specifications served as shortcuts for “vibe preferences” (based on “vibe coding”  (Roose, 2025)); where users did not full specify what they were looking for, but challenged the system to interpret rich input. Participants largely felt starting with an LM prompt felt like “a very natural way to get [the intention] down at first” (P6). P4 also conveys this elegance:
@@
P11, who had experience using other feed building tools, explained that in other interfaces they were “just basically writing Boolean operators about these combinations of words,” whereas Bonsai let them “just kind of stream of consciousness talk about the things you want to see, which is much more intuitive, right? It gets you closer to that— I know what I want to see, rather than trying to sit there and think of keywords.”
@@
On the other hand, participants also pointed out specific times they could struggle or even feel uncomfortable voicing their preferences in natural language. P10 noted this tension from the get-go: Bonsai “made me realize that it can be difficult to articulate in natural language our specific interests.” P8 struggled with defining a certain type of post they did not want to see: “because it was not an advertisement, it was just, like, someone saying, hey, this is a tool” and lacked “the general absence of a clear definition on what it was that I didn’t want to see.” When making edits, P8 stated “I think it was because I was not really sure about what I was even seeing, and then having communicate that [felt challenging].” P6 specifically tried to describe a feed “personalized to my specific humor” and found that “pretty hard to do.” One participant said they would prefer to see less of a certain type of political content that negatively affected their mental health, but felt that writing this desire out in words felt “wrong.”
@@
Due to these constraints with expressing preferences in natural language, participants sought to combine explicit and emergent preferences. For example, some participants mentioned that they would benefit from providing examples of posts they liked/disliked without having to describe them in natural language.
@@
“There’s that famous Supreme Court case where the justice said: ‘I can’t define porn, but I know it when I see it.’ [^1] From “I know it When I See It:” A History of the Definition of Obscentity: “I shall not today attempt further to define the kinds of material I understand to be embraced within that shorthand description [”hard-core pornography”], and perhaps I could never succeed in intelligibly doing so. But I know it when I see it, and the motion picture involved in this case is not that”. I think there’s a general sense, where I think even when I’m trying to describe what my current feeds are like, and then what my preferred ones are like… […] I find those very challenging to describe and put words to. I think it’s more of an experience.” -P6
5.2. Controlling explicit and personalized social media feeds

### v2

@@ 3.1. Design Goals
The central purpose of Bonsai is to realize a feed that is both intentional and personalized. We operationalize this purpose by pursuing three design goals detailed below: intent fidelity, ease of use, and procedural transparency.
@@
Explicitly stated intentions govern the sourcing, filtering, and ranking of content in the feed. Users should have an interface that allows them to specify directly what they want—and do not want—to see. This goal responds to prior work (§2.1) showing that optimizing for revealed preferences (e.g., clicks, dwell time) often drives regrettable or addictive use  (Tran et al., 2019b; Cho et al., 2021; Milli et al., 2025). This lets users isolate the personalization signal to their instructions, preventing hidden optimization toward engagement and ensuring that their social media use aligns with their intended goals.
3.1.2. Ease of use.
Users set and adjust feed goals without specialized vocabulary or interfaces. The interface between users and the feed algorithm should be simple, flexible, and expressive. Lowering the cost of articulation broadens who can personalize feeds and reduces drift toward opaque defaults. This design goal draws on HCI systems (§2.2) that show lightweight, accessible interventions are effective in helping people regulate their technology use (Hiniker et al., 2016; Grüning et al., 2023; Ko et al., 2015). Ease of use is also important to ensure intent fidelity, as increasing the flexibility and ease with which users specify their intentions ensures alignment between intention and feed.
@@
The feed creation process is procedurally transparent. Users should clearly understand the process by which their feed is generated, and how different aspects of the configuration impact the resulting feed. This design goal builds on research (§2.3) showing that people frequently misunderstand how feeds operate and consequently miscalibrate their trust or behavior (Eslami et al., 2016; Rader and Gray, 2015), as well as on scholarship critiquing “black-box” notions of transparency and urging attention to the processes through which systems operate (Ananny and Crawford, 2018). Procedural transparency is essential to making the system easy to use: only when users understand how their input shapes the system’s outputs can they refine and trust it.
@@ 3.2. Interface Design and System Walkthrough
We designed Bonsai based on the aforementioned landscape analysis and design goals. During prototyping, we iteratively refined defaults and interaction flows in response to formative feedback from project researchers and informal demos with colleagues. Two key changes from this process were: 1) the creation of an initial prompt interface (see §3.2.1) to bootstrap the feed creation process; and 2) the addition of different ranking algorithms that prioritize recency, popularity, or topical fit (see §3.2.4).
We implemented Bonsai for Bluesky for normative and pragmatic reasons. Bluesky is open, decentralized, built on the AT protocol, and designed to give users more control over their experience, content, and data. The AT protocol makes it easier to implement a feed curation system, and Bluesky users are also more likely to be interested in such an application. The user experience (Figure 3) unfolds through four steps, which we describe below.
@@ 3.2.2. Selecting sources
Next, users specify the sources that will populate their feed (Fig. 3B). Based on the initial description, Bonsai suggests feeds, starter packs, lists, search queries, accounts, and hashtags. Users can accept, edit, or remove these suggestions, and manually add new sources. Sources can be accounts, other feeds, search terms, and ‘starter packs’ (curated sets of accounts).
@@ 3.2.5. Generating and using the feed
After completing these steps, the user can click “Generate Now” to initiate the feed generation process. A few seconds after initiating generation, Bonsai may suggest additional sources to the user, based on the edits made to the content preferences. After generation, the feed appears as a custom feed option in the user’s Bluesky app. Bonsai continues to update the feed in the background. If users wish to edit their feed, they can always make changes on the feed customization page. They can modify preferences, add new exclusions, adjust source selections, and even create new feeds.
@@ 3.3. System Architecture
We propose a generalizable system architecture for generating intentional and personalized social media feeds. This architecture comprises four components (Planning, Sourcing, Curation, and Ranking), depicted in Fig. 4. While our implementation targets Bluesky, the modular design is platform-agnostic: each component can be adapted to different ecosystems by plugging in platform-specific APIs, content databases, or models. We chose Bluesky because its open, decentralized AT Protocol makes external feed customization feasible.
@@ 3.3.1. Planning
The Planning component translates a user’s natural language prompt into an initial feed configuration. This includes suggested sources (e.g., feeds, lists, hashtags) and content preferences (e.g., “include posts about” and “limit posts about”). Because this step involves semi-autonomous interpretation of natural language, we employ a language model (LLM) agent. In our Bluesky prototype, the agent queries an internal SQLite database of candidate sources (with previously scraped feeds, lists, and ‘starter packs’) and generates search queries and filter suggestions from the prompt. The resulting configuration is surfaced to the user for review and refinement. Our implementation used GPT-4o via Azure; the Planning stage can be instantiated with other models or retrieval systems, depending on the target platform. We provide full details of the feed, starter-pack, and list database construction in Appendix B, including our scraping procedure, filtering criteria, and the final counts (1,277 feeds, 205 starter packs, and 300 lists). We provide additional details about the planner agent in Appendix H.
@@ 3.3.3. Curating
The Curating component applies the user’s stated preferences to the sourced posts and assigns each item a relevance score. In principle, this stage can be implemented with classifiers or rules tailored to the platform; in our case, we leveraged a multimodal LM (GPT-4o) to handle Bluesky’s mix of text and images and to align directly with users’ natural language prompts. Each post received a score on a 0–10 scale: “strongly prefer” (8–10), “prefer” (5–7), “unspecified” (3–4). “show less often” (1–2), “never show” (0). Posts scored as zero were filtered out. Although our prototype relied on a relatively heavyweight model, the curation logic could be reimplemented with lighter models, heuristics, or hybrid pipelines, depending on deployment constraints. We provide the full prompts used in Appendix G.
@@ 3.3.4. Ranking
The Ranking component orders the curated posts into a final feed according to user-specified priorities. We adopted three widely used dimensions from existing feed algorithms: relevance, recency, and popularity (Covington et al., 2016; Guess et al., 2023). Relevance scores came from the Curating stage; popularity was computed as a weighted engagement score (likes + 3 $\times$ reposts + 2 $\times$ replies); and recency was calculated via reverse-chronological order. To combine these, we applied a Weighted Borda Count, with weights set by the user’s chosen ranking style (e.g., “Focused,” “Fresh,” “Balanced,” “Trending”). This structure generalizes readily: other platforms may substitute alternative engagement metrics, quality signals, or normative objectives. We provide more information and the implementation of the Weighted Borda Count algorithm in Appendix C.
@@ 4. Methods
We conducted a field study with 15 participants who used Bonsai for a median of 11 days (min: 6 days; max: 25 days; avg: 12.27 days). We aimed for approximately one week of use for each participant, though the exact number of days varied based on availability. We interviewed participants at the beginning and end of the study in order to understand the details of their experience. The study lasted one month, from July 29th to August 29th, 2025.
@@ 4.1. Participants
We recruited participants primarily through social media, creating a visual flyer that we shared on Bluesky and in relevant Slack groups. We also conducted targeted outreach on Bluesky to engage frequent users of the platform and to enlist non-academics in the study. From those who filled out the interest survey, we contacted participants we could validate as real people and as regular Bluesky users. In total, 15 participants enrolled in the study. Participants received a \$50 gift card for completing both interviews, and all procedures were approved by Princeton IRB (#18334). Throughout this paper, we anonymize participant names using PX nomenclature and further obscure highly specific stories or feed intentions that could be recognizable.
We took a broad stance on participant inclusion: the only requirement was that participants be 18 or older and actively use Bluesky. Despite this broad criterion, likely due to the skew of our own outreach, all participants except one were academics. We purposefully sought participants with different levels of Bluesky and custom feed usage, and five participants already used other custom feeds on Bluesky. Eleven participants lived in the US, two in the UK, and the others in Canada, Austria, and Spain.
@@ 4.2. Procedure
The study comprised an onboarding interview, at least a week of using Bonsai feeds, and a follow-up interview. Both interviews were semi-structured and lasted 45 minutes. In the first interview, we explored participants’ general social media habits, asking what value they currently derive from feeds, what they feel is missing, and what they wish feeds would provide. We then walked them through the feed generation process in Bonsai, asking them to think aloud as they selected sources, wrote prompts, and configured ranking options. During the trial week, participants were encouraged to use Bonsai in their everyday browsing, creating and refining feeds as they wished.
The second interview took a more reflective and evaluative form. We asked participants to describe how they attempted to convey their intent in the feed, whether they felt successful, and what challenges they encountered. We also probed the broader pros and cons of intentionality, natural language input, and feed-level control, as well as participants’ overall impressions of the system. In several cases, we revisited the feed creation process to observe how their strategies evolved. In both interviews, we examined the feeds participants had generated, discussing whether the content matched their expectations and probing moments when posts seemed out of place. We include the full interview protocols in the Appendix sections E and F.
@@ 4.3. Data Analysis
Participant interviews were recorded and automatically transcribed on Zoom. We uploaded the anonymized transcripts to Atlas.ti, where they were further refined for accuracy. The first and second authors then engaged in open coding for two full interviews each, discussing and honing the initial codes with the broader research team. As additional interviews were being conducted, we iteratively expanded the codebook and discussed emerging themes. Some of the most relevant initial higher-level themes we identified included social media motivations and goals, relationship between control and agency, describing intent in natural language, and feed curation effort. Each of these had several subcodes, for example, describing intent in natural language encompassed preferences hard to capture in natural language, perceptions and folk theories of LLMs, intuitive, powerful, and other. This codebook and discussions with the research team guided interviews, for example, focusing more on effort and transparency since these were yielding compelling insights from participants. We continued to iterate on these themes through discussions as interviews were conducted, eventually uniting them into three main overarching sections which align with our design goals: imprinting intentions on social media, controlling intentional and personalized social media feeds, procedural transparency of intentional and personalized feeds.
In all, we conducted fifteen interviews. Interviews continued until thematic saturation was reached (no new codes added after two interviews) after fourteen interviews. We considered this sample size sufficient given prior studies identified that the mode number of participants in HCI studies is twelve (Caine, 2016). Nonetheless, we conducted one additional interview to capture more non-academic perspectives. We note that, given our small and predominantly academic sample of active Bluesky users, our findings are not intended to be statistically generalizable, but to characterize how this particular population appropriates and experiences LM-mediated intentional feeds.
To add richness to our findings, we also captured the Bonsai interaction logs and the natural language prompts from participants. The natural language prompts were tagged with specific user requests, such as a topical focus or a request to filter specific content. Additionally, we manually analyzed the logs for each participant to identify the actions they took on the platform between interviews, such as editing a feed or creating a new feed. These results are incorporated throughout the findings section.
To ensure that Bonsai-generated feeds were adequate, we had 7 annotators each annotate 60 posts. For each post, annotators were given user-defined prompts and asked to judge whether posts were ”perfectly relevant,” ”relevant,” ”not relevant,” ”should have been excluded,” or ”not sure.” Each annotator annotated posts for two different feeds (30 each). 70% of posts were rated as relevant or better (26% ”perfectly relevant,” 44% ”relevant”). Further, 16% were rated ”not relevant,” and 7% ”should have been excluded.” We consider these annotations evidence that the system can generate feeds that largely align with users’ stated intentions, especially given that the 14 considered feeds were randomly selected (and not necessarily the best-performing ones). We had an eighth annotator label posts that had been previously annotated, finding substantial inter-annotator agreement ($\kappa=0.77$).
@@ 5. Results
We discuss how participants used Bonsai during the study period, focusing on the prompts they authored, the actions they performed, and the ways they evaluated their feeds. We structure our findings around our three design goals: (1) how participants imprinted their intentions on social media feeds, (2) how they steered and maintained intentional and personalized feeds, and (3) how they made sense of the procedural transparency of feed generation. Our 15 participants built between 1 and 4 feeds each for a total of 36 feeds. They manually re-generated their feeds between 2 and 16 times for a total of 103 times. Per week, they consumed Bonsai feeds between 8 and 264 times each for a total of 1413 requests.
@@ 5.1. Imprinting intentions on social media feeds
5.1.1. Participants specified topical interests using Bonsai to create both general and specialized feeds.
@@
In the remaining 16 feeds, participants used Bonsai to build feeds for specific topics or hobbies they were interested in. P7 created a feed focused on cinematography (“I want to see film festival news, film critics, artists’ sharing their experiences or new releases, experimental / AI films”), P11 on college sports (“I want to see all posts related to US college basketball”), and P15 (“a videogame development feed.”) Before using Bonsai, some participants had already used complementary topical feeds, such as P6, who followed a feed of political science journals, while P9 would repeatedly search for common keywords by clicking on the same keywords from their saved search history – effectively creating a pseudo-feed. For this participant, Bonsai offered a much easier way to create a topical feed suited to their interests.
Regardless of whether participants wanted a feed with one or multiple topics, they often leveraged the flexible LLM-driven nature of Bonsai to specify complex sub-topics or go beyond surface-level feed topics. In this vein, P15 created a feed that was specifically about video game development and building, rather than a more general video game feed; while P7 wanted to see news about experimental/AI films as opposed to commercial movies. For participants who were already relying on topic-based feeds, Bonsai offered a way to hone in on specific sub-parts of a topic. For example, P11 followed sports feeds already but wanted some college basketball teams to be featured more heavily than others. These kinds of within-topic customizations had been hard for participants to achieve before Bonsai: “this [feed] was really, really good, this is exactly what I wanted […] it was a lot easier to dial this one in” (P11).
@@
These desires for better filters stemmed from participants’ negative social media experience, particularly on Twitter. Some users sought to improve their mental health on platforms like X, which P8 called “toxic” and “unbearable,” P10 reported feeling “anxious when using it,” and P6 felt made them “deeply unhappy.” Others felt that engagement-based feeds simply were not effective at filtering out content, even when given explicit instructions. P8 said that on Twitter they “would always say, okay, I want to see less of that, and I wouldn’t see less of it,” while P15 held the same opinion of Bluesky, where “even though you tell it ‘show less of this’, it keeps displaying it.” P4 told the story of a user whose academic content they enjoyed, but who also posted frequently about football—a topic they had no interest in. In response, P4 tried to block football-relevant keywords, but found “with the existing methods, you can achieve certain results, they are certainly useful in some contexts, but they’re not very specific, they’re not very, sort of fine-grained.” P13 shared a similar experience, saying ‘on Bluesky, I would say [filtering] is terrible. I have set up 30 keywords to mute things, so that I don’t see political content on Bluesky, but it still manages to creep onto my feed.”
@@
Although participants felt intentional feeds could be liberating, some participants felt perfectly in control of their existing engagement-based feeds. P13 for example explained “on Twitter, […] I could just hit ’not interested,’ and then I won’t see stuff,” and that they feel “I have a lot of self-control on what I click on to shape my algorithmic feed, and so because of that, I can get a really good feed on Twitter.” P9 similarly felt confident that platforms could infer their intent correctly: “once I give somebody a thumb [or a] heart, [the platform] can more quickly capture my interests and then start recommending more things similar to that”. Even those who felt more lukewarm about their ability to manage engagement-based feeds still exerted additional controls over their feeds by blocking and muting accounts and keywords to filter out content they did not wish to see.
@@
Through our design, participants cleverly used natural language to convey specific and nuanced desires. For example, participants used humor and specific cultural references to help the AI “situate” itself within the type of desired content. Nuanced references could allow participants to convey complex semantic ideas with just a few words. These specifications served as shortcuts for “vibe preferences” (based on “vibe coding”  (Roose, 2025)); where users did not fully specify what they were looking for, but challenged the system to interpret rich input. Participants largely felt starting with an LM prompt felt like “a very natural way to get [the intention] down at first” (P6). P4 also conveys this elegance:
@@
On the other hand, participants also pointed out specific times they could struggle or even feel uncomfortable voicing their preferences in natural language. P10 noted this tension from the get-go: Bonsai “made me realize that it can be difficult to articulate in natural language our specific interests.” P8 struggled with defining a certain type of post they did not want to see: “because it was not an advertisement, it was just, like, someone saying, hey, this is a tool” and lacked “the general absence of a clear definition on what it was that I didn’t want to see.” When making edits, P8 stated “I think it was because I was not really sure about what I was even seeing, and then having to communicate that [felt challenging].” P6 specifically tried to describe a feed “personalized to my specific humor” and found that “pretty hard to do.” One participant said they would prefer to see less of a specific political conflict that negatively affected their mental health, but felt that writing this desire out in words felt “wrong,” since others had to experience the conflict and could not tune it out. P1 also felt it was ethically complex to intentionally define the types of news they want to see, saying that although they are “always creating an echo chamber [in implicit feeds], it just feels bad to explicitly write these rules.”
@@
Due to these constraints in expressing preferences in natural language, participants sought to combine intentional and emergent preferences. For example, some participants mentioned that they would benefit from providing examples of posts they liked/disliked without having to describe them in natural language.
@@
“There’s that famous Supreme Court case where the justice said: ‘I can’t define porn, but I know it when I see it.’ [^2] From “I know it When I See It:” A History of the Definition of Obscenity: “I shall not today attempt further to define the kinds of material I understand to be embraced within that shorthand description [”hard-core pornography”], and perhaps I could never succeed in intelligibly doing so. But I know it when I see it, and the motion picture involved in this case is not that”. I think there’s a general sense, where I think even when I’m trying to describe what my current feeds are like, and then what my preferred ones are like… […] I find those very challenging to describe and put words to. I think it’s more of an experience.” -P5
5.2. Steering social media feeds